OpenStudy (anonymous):

Write the equation of the line in slope-intercept form with a slope of 3 and a y-intercept of -2.

OpenStudy (anonymous):

pls help me

OpenStudy (anonymous):

So slope intercept form is y=mx+b correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So the first step is to insert the numbers into the equation. m is the slope and y is the y-intercept

OpenStudy (anonymous):

so -2=3x+b

OpenStudy (anonymous):

So it's -2=3x+b

OpenStudy (anonymous):

Did they give you a number for x?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

so thats why i am confused

OpenStudy (anonymous):

wait so do i just put -2=3x+b

OpenStudy (anonymous):

No, you need b. But I don't know how to find that.

OpenStudy (anonymous):

oh ok can you help me find someone that can?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@JA1

OpenStudy (anonymous):

So sorry I was no help!

OpenStudy (anonymous):

its ok @SoccerManiac

jimthompson5910 (jim_thompson5910):

In general y = mx+b is slope intercept form m is the slope b is the y-intercept so because the slope is 3, this means m = 3 because the y-intercept is -2, we know that b = -2 plug this in to get y = mx + b y = 3x + (-2) y = 3x - 2

OpenStudy (anonymous):

wait so b is -2?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

ok so basically there is no y? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

what do you mean?

OpenStudy (anonymous):

like in y=mx+b @jim_thompson5910

jimthompson5910 (jim_thompson5910):

well x and y are variables that are allowed to change (and aren't fixed) so that's why you can't just replace y with some fixed number like you can with m and b

OpenStudy (anonymous):

so the answer would be y=3x-2?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

yep, that is correct

OpenStudy (anonymous):

@jim_thompson5910 ok well thank u so much

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (dan815):

ocean are you good with this stuff?

OpenStudy (dan815):

you should check out youtube khan academy videos about this stuff, u ll learn a lot of stuff